when n is divided by 5 the remainder is 2 what is the remainder when square of n is divided by 5?
Answer by stanbon(75887)

when n is divided by 5 the remainder is 2 what is the remainder when square of n is divided by 5?
------
n = 5k + 2 where k is an integer
-----
n^2 = 25k^2+20k+4
----
Divide that by 5 to get:
n^2 = 5(5k^2+4k)+ 4/5
-------
Remainder = 4/5
=====================
Cheers,
Stan H.
=====================